1.The pdf of a random variable x is given by f(x)={k(1-x),0<x<1 both inclusive

{0, elsewhere

Find the value of the constant k,the cdf of x and the value of m such that G(x)=1/2

2.Find the cdf of a random variable y where the pdf is given by

(a)f(x)={1/3,0<X<1 both inclusive

{1/3,2<x<4 both inclusive

{0,elsewhere

(b)f(x)={x/2,0<x<1 both inclusive

{1/2,1<x<2 both inclusive

{(3-x)/2,2<x<3 both inclusive

3.If the cdf of a random variable y is given by F(X)=1-9/y2 for y>or equal to3

and F(x)=0 for y<3 find the P(X<or equal to 5),p(x>8) and the pdf of x.



1
Expert's answer
2021-07-12T13:18:44-0400

1. "1=\int f(t)dt=\int\limits_{0}^1k(1-t)dt=k(t-t^2/2)|_0^1=k/2"

Therefore, k=2.

The CDF (cumulative distribution function) is "F(x) =\int\limits_{-\infty}^x f(t)dt".

"F(x)=0", if "x<0";

"F(x)=1", if "x>1";

"F(x)=\int\limits_{0}^x 2(1-t)dt=(2t-t^2)|_0^x=2x-x^2", if "x\in[0,1]".

The median of distribution is a such x, that "F(x)=1\/2".

"2x-x^2=1\/2"

"(x-1)^2=1\/2"

"x=x_1=1-1/\sqrt{2}" , since the other root "x_2=1+1/\sqrt{2}>1" and "F(x_2)=1"


2. (a) "f(x)=1\/3", if "X\in[0,1]\cup [2,4]" and "f(x)=0" elsewhere.

The CDF is "F(x) =\int\limits_{-\infty}^x f(t)dt".

"F(x)=0", if "x\leq 0";

"F(x)=\int\limits_{0}^{x}\frac{1}{3} dt=x/3", if "x\in[0,1]";

"F(x)=1/3+\int\limits_{0}^{x}0 dt=1/3", if "x\in[1,2]";

"F(x)=\frac{1}{3}+\int\limits_{2}^{x}\frac{1}{3} dt=\frac{1}{3}+\frac{x-2}{3}=\frac{x+1}{3}", if "x\in[2,4]";

"F(x)=1", if "x>4".

(b) Given the pdf "f(x)=x\/2" , if "0\leq x\leq1"

"f(x)=1\/2" , if "1\leq x\leq2"

"f(x)=(3-x)\/2" if "2\leq x\leq 3"

is not a pdf of a random variable, since

"\int f(t)dt=\int\limits_{0}^{1}\frac{t}{2}dt+\int\limits_{1}^{2}\frac{1}{2}dt+\int\limits_{2}^{3}\frac{3-t}{2}dt=\frac{1}{4}+\frac{1}{2}+\frac{1}{4}"

The CDF (cumulative distribution function) is "F(x) =\int\limits_{-\infty}^x f(t)dt".

"F(x)=0", if "x\leq 0";

"F(x)=\int\limits_{0}^{x}\frac{t}{2} dt=x^2/4", if "x\in[0,1]";

"F(x)=\frac{1}{4}+\int\limits_{1}^{x}\frac{1}{2} dt=\frac{1}{4}+\frac{x-1}{2}=\frac{2x-1}{4}", if "x\in[1,2]";

"F(x)=\frac{3}{4}+\int\limits_{2}^{x}\frac{3-t}{2} dt=\frac{3}{4}+\frac{1}{2}(x-2)-\frac{(x-2)^2}{4}=\frac{-x^2+6x-5}{4}", if "x\in[2,3]";

"F(x)=1", if "x>3".


3. If the cdf of a random variable Y is given by "F(y)=1-9\/y^2" for "y\geq 3" and "F(x)=0" for "y<3", then "P(Y<5)=F(5)=1-9\/5^2=0.64".

The PDF equals to "f(y)=F'(y)". Hence

"f(y)=18\/y^3", for "y\geq 3"

"f(y)=0", for any other value of y.